# Using MultiPrecisonArrays.jl: Iterative Refinement in Julia

### v7

@@ 5.1 Terminating the while loop
[…]
because we have seen several instances where IR can waste a few iterations and make very little progress. Our value of $r_{max}$ trades a bit of residual reduction for time.
In this paper we count iterations as residual computations. This means that the minimum number of iterations will be two. Since we begin with ${\bf x}=0$ and ${\bf r}={\bf b}$, the first iteration computes ${\bf d}={\bf U}^{-1}{\bf L}^{-1}{\bf b}$ and then ${\bf x}\leftarrow{\bf x}+{\bf d}$, so the first iteration is the output of a low precision solve. We will need at most one more iteration to test for a meaningful residual reduction.
@@ 5.2 Is $O(N^{2})$ really negligible
In this sectiion TR = TW = Float64 and TF = TS = Float32, which means that the iterprecision transfers in the triangular solvers are done in-place. We terminate on small residuals.
The premise behind IR is that reducing the $O(N^{3})$ cost of the factorization will make the solve faster because everything else is $O(N^{2})$ work. It’s worth looking to this.
We will use the old-fashioned defintion of a FLOP as an add, a multiply and a bit of address computation. So we have $N^{2}$ flops for any of (1) matrix-vector multiply ${\bf A}{\bf x}$, (2) the two triangular solves with the LU factors $({\bf L}{\bf U})^{-1}{\bf b}$, and (3) computaion of the $\ell^{1}$ or $\ell^{\infty}$ matrix operator norms $\|{\bf A}\|_{1,\infty}$.
@@
A single IR iteration costs a matrix-vector product in precision TR and a triangular solve in precision TF for a total of $23N^{2}/$ 2 TR-FLOPS. Hence a linear solve with IR that needs $n_{I}$ iterations costs
@@
[…]
@@ 5.3 Interprecision Transfers: Part I
The meaning of ${\bf d}={\bf U}^{-1}{\bf L}^{-1}{\bf r}$ is more subtle. The problem is that the factors ${\bf U}$ and ${\bf L}$ are store in low precision and ${\bf r}$ is a high precision vector. LAPACK will convert ${\bf L}$ and ${\bf U}$ to the higher precision ‘‘on the fly’’ with each mixed precision binary operation at a cost of $O(N^{2})$ interprecision transfers. The best way to understand this is to recall that if $TH$ is a higher precision than $TL$ $a,b\in{\cal F}_{H}$ and $c\in{\cal F}_{L}$ that
$fl_{H}(a*c+b)=fl_{H}(a+I_{L}^{H}(c)+b).$
For now, $TH=TW$ and $TL=TF$, but we need to think of the general case in § 6. As we will see this interprecision transfer can have a meaningful cost even though the factorization will dominate with $O(N^{3})$ work.
@@ 5.4 Convergence Theory
To estimate the first term we look at the component-wise backward error [7]. If $13Nu_{F}<$ 1 then
(9) $|{\hat{\bf L}}{\hat{\bf U}}-I_{W}^{L}{\bf A}|\leq\gamma_{3N}(u_{F})|{\hat{\bf L}}||{\hat{\bf U}}|.$
@@
(10) $\begin{array}[]{ll}\|{\bf M}_{IR}\|&\leq u_{F}\|{\bf A}\|+\gamma_{k}(u_{F})\|{\hat{\bf U}}^{-1}{\hat{\bf L}}^{-1}\|\|{\hat{\bf L}}\|\|{\hat{\bf U}}\|\\ \\ &=u_{F}\|{\bf A}\|+\gamma_{k}(u_{F})\kappa({\hat{\bf L}})\kappa(\hat{\bf U}).\end{array}$
@@
(12) $\begin{array}[]{ll}\|{\bf M}\|&\leq u_{F}(\|{\bf A}\|+\|{\hat{\bf U}}^{-1}{\hat{\bf L}}^{-1}\|13\sqrt{N}\|{\hat{\bf L}}\|\|{\hat{\bf U}}\|)\\ \\ &\leq u_{l}(\|{\bf A}\|+13\sqrt{N}\kappa({\hat{\bf L}})\kappa(\hat{\bf U})).\end{array}$
@@ 5.4.2 Limiting Behavior of IR
(15) $\alpha=u_{F}(\|{\bf A}\|+13\sqrt{N}\kappa({\hat{\bf L}})\kappa(\hat{\bf U})).$
@@
As is standard, when one computes a residual ${\bf r}$ the computed value ${\hat{\bf r}}$ has an error [7]
@@
We will analyze the progress of IR while only considering the errors in the the residual computation. So we compute
@@
When we terminate the iteration when $\|{\bf r}\|/\|{\bf b}\|$ is small we are ignoring the $\|{\bf A}\|\|{\bf x}^{*}\|$ term in $\xi$, which one reason we must take watch for stagnation in our solver.
6 Evaluating the residual in extended precision
This idea comes from [20] and I am using the notation from [6, 1]. The idea is to evaluate the residual in a precision TR higher than the working precision. If you do this, you should store both the solution and the residual in precision TR and to interprecision transfers on the fly. In that case you are really solving a promoted problem [15]
$(I_{W}^{R}A)x=I_{W}^{R}b$
and, by driving the residual to a small value can mitigate ill-conditioning to some degree. MultiPrecisionArrays.jl allows you to do this with the multiprecision factorization you get from mplu. I may add this feature to the Krylov-IR solvers later.
The classic example is to let TR and TF be single precision and TR be double. The storage penalty is that you must store two copies of $A$, one for the residual computation and the other for the factorization. Here is an example with a badly conditioned matrix. You must tell mplu to factor in the working precision and then use the kwargs in the solver to set TR.
julia> using MultiPrecisionArrays
julia> using MultiPrecisionArrays.Examples
julia> N=4096; alpha=799.0; AD=I - alpha*Gmat(N);
# conditioning is bad
julia> cond(AD,Inf)
2.35899e+05
# Set up the single precision computation
julia> A = Float32.(AD); xe=ones(Float32,N); b=A*xe;
# Make sure TF is what it needs to be for this example
julia> AF = mplu(A; TF=Float32);
# Use the multiprecision array to solve the problem, set TR.
julia> mrout = \(AF, b; reporting=true, TR=Float64);
# look at the residual history
julia> mrout.rhist
5-element Vector{Float64}:
9.88750e+01
1.67735e-05
9.23976e-10
9.37916e-13
9.09495e-13
# Compare results with LU and the exact(?) solution
[…]
julia> [norm(b - A*xr, Inf) norm(b - A*xs, Inf)]
[…]
1.29700e-04 1.41144e-03
# So the residual is better. What about the difference from xe?
julia> [norm(xr - xe, Inf) norm(xs - xe, Inf)]
[…]
8.47816e-04 8.82089e-04
# Nothing exciting here. You have to wonder what this all means.
# Finally, how did we do with the promoted problem?
julia> AP=Float64.(A);
julia> norm(b - AP*mrout.sol, Inf)
7.10543e-13
# Which is what I said it was above.
So, is the solution to the promoted problem better than the exact solution I used to build the problem?
AF = mplu(A; TF=Float32);
with
AF = mplu(A);
The iteration fails to converge because $AF$ is not accurate enough.o
julia> AF = mplu(A; TR=Float64);
julia> mout16=\(AF, b; reporting=true);
julia> mout16.rhist
5-element Vector{Float64}:
9.88750e+01
3.92614e+00
3.34301e-01
2.01975e-01
2.24576e-01
6.1 IR-Krylov
MultiPrecisionArrays.jl supports IR-GMRES and IR-BiCGSTAB for TR > TW. You get this to work just like in mplu by using the keyword argument TR. We will continue with the example in this section and do that.
julia> GF = mpglu(A; TR=Float64);
julia> moutG = \(GF, b; reporting=true);
julia> moutG.rhist
6-element Vector{Float64}:
9.88750e+01
2.23211e-04
9.61252e-09
1.26477e-12
8.10019e-13
8.66862e-13
# You need several iterations because the default is 10 Krylov vectors
# And we got the solution to the promoted problem...
[…]
julia> norm(xp-moutG.sol, Inf)
6.52844e-12
# IR-BiCGSTAB should take fewer iterations because there's no storage
julia> BF = mpblu(A; TR=Float64);
julia> moutB = \(BF, b; reporting=true);
julia> moutB.rhist
4-element Vector{Float64}:
9.88750e+01
2.16858e-11
8.38440e-13
8.81073e-13
julia> norm(xp - moutB.sol, Inf)
1.36534e-11
@@ A.1 Double-Single IR
[…]

### v8

@@ 5.1.2 Protection Against Stagnation
even if the small residual condition is not satisfied. The recommendation in [8] is $R_{max}=.5$ and that is the default in MultiPrecisionArrays.jl .
The paper [8] also recommends that on put a limit litmax = 5 on the number of IR iterations. We use a higher value of litmax = 10 in MultiPrecisionArrays.jl .
In this paper we count iterations as residual computations. This means that the minimum number of iterations will be two. Since we begin with ${\bf x}=0$ and ${\bf r}={\bf b}$, the first iteration (iteration 0) computes ${\bf d}={\bf U}^{-1}{\bf L}^{-1}{\bf b}$ and then ${\bf x}\leftarrow{\bf x}+{\bf d}$, so the first iteration is the output of a low precision solve. We will need at most one more iteration to test for a meaningful residual reduction. Iterations after iteration 0 cost more because they do a meaningful residual computation.
5.1.3 Changing the parameters
The parameters and defaults are
  • $C_{r}=1.0$
  • $C_{e}=1.0$
  • $R_{max}=.5$
  • $litmax=10$
We store the parameters in a TERM structure, which define in the main file for MultiPrecisionArrays.jl .
struct TERM
Cr::Real
Ce::Real
Rmax::Real
litmax::Int
end
We create one TERM structure for the defaults term_parms_default. The solvers take a TERM structure as a kwarg, with term_parms_default as the default.
To change the parameters use the update_parms function. This function makes a TERM structure for you to pass to the solvers.
function update_parms(; Cr=Cr_default, Ce=Ce_default,
Rmax=Rmax_default, litmax=litmax_default)
term_parms=TERM(Cr, Ce, Rmax, litmax)
return term_parms
end
Here is an example with the ill-conditioned problem we’ve been using in other examples. In this example we decrease Rmax and see that the solve takes fewer iterations with no change in the residual quality.
julia> using MultiPrecisionArrays
julia> using MultiPrecisionArrays.Examples
julia> N=512; A=I - 799.0*Gmat(N); AF=mplu(A);
julia> b=ones(N);
julia> mout=\(AF,b; reporting=true);
# Termination on a residual norm increase.
julia> mout.rhist
6-element Vector{Float64}:
1.00000e+00
4.39096e-03
2.85170e-07
4.30167e-11
6.05982e-12
6.34648e-12
julia> term_parms=update_parms(;Rmax=.1);
julia> mout2=\(AF,b; reporting=true, term_parms=term_parms);
# Termination on minimal decrease in the residual norm.
julia> mout2.rhist
5-element Vector{Float64}:
1.00000e+00
4.39096e-03
2.85170e-07
4.30167e-11
6.05982e-12
@@ 5.2 Is $O(N^{2})$ really negligible
In this section TR = TW = Float64 and TF = TS = Float32, which means that the iterprecision transfers in the triangular solvers are done in-place. We terminate on small residuals.
The premise behind IR is that reducing the $O(N^{3})$ cost of the factorization will make the solve faster because everything else is $O(N^{2})$ work. It’s worth looking into this.
We will use the old-fashioned definition of a FLOP as an add, a multiply and a bit of address computation. So we have $N^{2}$ flops for any of (1) matrix-vector multiply ${\bf A}{\bf x}$, (2) the two triangular solves with the LU factors $({\bf L}{\bf U})^{-1}{\bf b}$, and (3) computation of the $\ell^{1}$ or $\ell^{\infty}$ matrix operator norms $\|{\bf A}\|_{1,\infty}$.
@@
A single IR iteration costs a matrix-vector product in precision TR and a triangular solve in precision TF for a total of $3N^{2}/2$ TR-FLOPS. Hence a linear solve with IR that needs $n_{I}$ iterations costs
@@
IR will clearly be better for large values of $N$. How large is that? Table 2 and Table 3 compare the time for factorization nd solve using $lu!$ and ldivldiv (cols 2-4) with the equivalent multiprecision commands mplumplu and $\backslash$. The final column is the time for computing $\|{\bf A}\|_{1}$, which would be needed to terminate on small normwise backward errors.
@@ 5.3 Interprecision Transfers: Part I
The meaning of ${\bf d}={\bf U}^{-1}{\bf L}^{-1}{\bf r}$ is more subtle. The problem is that the factors $\bf U$} and $\bf L$} are store in low precision and $\bf r$} is a high precision vector. LAPACK will convert $\bf L$} and $\bf U$} to the higher precision ‘‘on the fly’’ with each mixed precision binary operation at a cost of $O(N^{2})$ interprecision transfers. The best way to understand this is from (1), which implies that if $\bf L$} and $\bf U$} are in precision $TF$ and $\bf b$} is in precision $TW$, then
$({\bf L}{\bf U})^{-1}{\bf b}=((I_{F}^{W}{\bf L})(I_{F}^{2}\mu))^{-1}{\bf b}.$
As we will see this interprecision transfer can have a meaningful cost even though the factorization will dominate with $O(N^{3})$ work.
@@ 5.4 Convergence Theory
To estimate the first term we look at the component-wise backward error [7]. If $3Nu_{F}<1$ then
[…]
@@
[…]
@@
[…]
@@ 5.4.2 Limiting Behavior of IR
(18) $).\alpha=u_{F}(\|{\bf A}\|+13\sqrt{N}\kappa({\hat{\bf L}})\kappa(\hat{\bf U}$)).
@@
As is standard, when one computes a residual $\bf r$} the computed value $^{\hat{\bf r}}$ has an error [7]
@@
We will analyze the progress of IR hile only considering the errors in the the residual computation. So we compute
@@
When we terminate the iteration when $\|{\bf r}\|/\|{\bf b}\|$ is small we are ignoring the $\|{\bf A}\|\|{\bf x}^{*}\|$ term in $\xi$, which is one reason we must take watch for stagnation in our solver.
5.4.3 Extended precision residuals
We consider the case $TR>TW$ and do interprecision transfers in the triangular solves on the fly. In the language of [4] the solver precision and working precision are the same.
The convergence result for this case (see Corollary 3.3, page A824, of [4] or also see [18]) is (stated very roughly)
Theorem 5.1.
Assume that $\|M_{IR}\|$ is sufficiently <1<1. Then the error in IR will be reduced until, at stagnation,
(21) $\frac{\|{\bf x}-{\bf x}^{*}\|_{\infty}}{\|{\bf x}^{*}\|_{\infty}}\leq 4(N+1)\kappa_{\infty}({\bf A})u_{r}+u_{w}.$
So, if $u_{r}\leq u_{w}^{2}$ and $(N+1)\kappa_{\infty}({\bf A})=O(u_{w}^{-1})$ then
(22) $\frac{\|{\bf x}-{\bf x}^{*}\|_{\infty}}{\|{\bf x}^{*}\|_{\infty}}=O(u_{w}),$
which is what we observed in the examples.
@@ A.1 Double-Single IR
Finally we mention a subtle programming issue. We made Table 4 with the standard commands for matrix-vector multiply (${\bf A}*{\bf x}$), factorization lu, and used $\backslash$ for the solve. Julia also offers non-allocating versions of these functions. In Table 5 we show how using those commands changes the results. We used mul! for matrix-vector multiply, lu! for the factorization, and ldiv! for the solve.